Nicolas Enterprises sells a product for $59 per unit. The variable cost is $30 per unit, while fixed costs are $114,376.

Determine (a) the break-even point in sales units and (b) the break-even point if the selling price were increased to $64 per unit.

a. Break-even point in sales units units
b. Break-even point if the selling price were increased to $64 per unit units
Answer #1

a.Contribution margin=Sales-Variable cost

=(59-30)=$29 per unit

Hence breakeven=Fixed cost/Contribution margin

=114,376/29=3944 units

b.Contribution margin=Sales-Variable cost

=(64-30)=$34 per unit

Hence breakeven=Fixed cost/Contribution margin

=114,376/34=3364 units

Break-even point in sales units 3944 units
Break-even point if the selling price were increased to $64 per unit 3364 units